OpenStudy (thunder33):

A total of 25 students have entered a spelling contest. There are 6 medals for first through sixth place that will be awarded. In how many different ways can the medals be awarded?

OpenStudy (kropot72):

This can be solved by treating it as the number of permutations of the 25 students, taken 6 at a time. Do you know the formula to use?

OpenStudy (thunder33):

no

OpenStudy (kropot72):

\[\large nPr=\frac{n!}{(n-r)!}\] Substituting the values given in your question, we get \[\large Number\ of\ ways=\frac{25!}{(25-6)!}=?\]

OpenStudy (kropot72):

Which reduces to: \[\large 25\times24\times23\times22\times21\times20\]